Andy has 8 coins made up of quarters and half dollars, and their total value is $2.50. How many quarters does he have?
Answer by stanbon(75887)

Andy has 8 coins made up of quarters and half dollars, and their total value is $2.50. How many quarters does he have?
---------------
2-variable solution:
Quantity: q + h = 8
Value:::25q+50h = 250 cents
----------------------------------
Multiply thru Quantity by 25 to get:
25q + 25h = 25*8
---
Subtract that equation from Value and solve for "h":
25h = 250-200
h = 2 (# of half-dollar coins)
---
Substitute into Quantity and solve for "q":
q + 2 = 8
q = 6 (# of quarter coins)
===============================
Cheers,
Stan H.